# Study of gravitational waves from phase transitions in three-component dark matter

### v2

@@ 4 Results
In this section, we calculate $\alpha$ as the strength parameter and $\beta$ as the inverse duration of the phase transition, for selected points, as shown in Table 4.
Table 4: Phase transition parameters based on (3.2) for the points in Table 2.
Point $T_{n}$(GeV) $\alpha$ $\frac{\beta}{H_{*}}$
A 116 0.1119 1004
B 101 0.0970 1469
Now by using the previous section’s relations and the results of table 4, we can determine the total background gravitational wave density from the electroweak phase transition as follows:
$\Omega_{\text{total}}(f)h^{2}\simeq\Omega_{\text{coll}}(f)h^{2}+\Omega_{\text{sw}}(f)h^{2}+\Omega_{\text{turb}}(f)h^{2}.$ (4.1)
Finally, figure 5 presents the total background gravitational wave density (4.1) versus GW frequency, utilizing data from table 4 for selected points. In figure 5, there are two curves for Lisa and BBO, which show the power of space-based laser detectors that have not yet been launched into space. As shown in figure 5, the Lisa detector does not have the capability to detect the background gravitational waves resulting from the electroweak phase transition for parameter space of B point; however, it is possible that other points like A, in our model exist where Lisa could detect the background gravitational waves from their electroweak phase transitions.
Refer to caption     Refer to caption
Figure 5: GW spectrum as a function of backgound gravitational frequency for points in Table 4.
In contrast, the BBO detector can detect a portion of these background gravitational waves for both points.
The background gravitational wave density from sound wave sources in both graphs in figure 5 is the main contributor to the total background gravitational wave. The maximum values in both graphs, indicate that the optimal GW frequency range for detecting background gravitational waves is between 0.003 and 0.5 Hz, with the best GW frequency at 0.02 Hz.
@@ 5 Conclusion
This paper introduces a model of dark matter in which particles exhibit three spin types: a vector particle, a scalar, and a spinor. In our model, the Higgs particle acts as a portal between Standard Model particles and dark matter. We identified parameter space points that exceed the detection capability of PandaX-4T and XENONnT, while their relic density matches the Planck satellite’s findings. We examined the electroweak phase transition for these points and showed that a potential barrier between the two minima of the effective potential at nucleation temperature can generate gravitational waves. The three parameters, $\alpha$, $\frac{\beta}{H_{*}}$, and nucleation temperature ($T_{n}$), are sufficient for gravitational wave calculations, independent of other model parameters. Our results indicate that the sensitivity of the LISA detector is insufficient for detecting cosmic background gravitational waves throughout the entire parameter space within the three-component dark matter model, whereas the BBO detector has the capability to confirm the existence of these waves.

### v3

@@ 4 Results
In this section, by using the previous section’s relations and the results of table 4, we can determine the total background gravitational wave density from the electroweak phase transition as follows:
$\Omega_{\text{total}}(f)h^{2}\simeq\Omega_{\text{coll}}^{env}(f)h^{2}+\Omega_{\text{sw}}(f)h^{2}+\Omega_{\text{turb}}(f)h^{2}.$ (4.1)
The values of $\alpha$ as the strength parameter and $\frac{\beta}{H_{*}}$ as the inverse duration of the phase transition, for selected points, as shown in Table 4.
Table 4: The background gravitational wave parameters resulting from the phase transition, obtained using (3.2), correspond to the points listed in Table 2.
Point $\alpha$ $\frac{\beta}{H_{*}}$ ${\Omega h^{2}}_{max}^{(total)}$ $f_{{\Omega h^{2}}_{max}^{(total)}}\;(HZ)$
A 0.123291 138.71 $8.63\times 10^{-13}$ $1.42\times 10^{-4}$
B 0.104421 1008.06 $1.06\times 10^{-14}$ $8.71\times 10^{-4}$
C 0.104247 1903.97 $2.95\times 10^{-15}$ $1.77\times 10^{-3}$
D 0.0386934 676.682 $1.51\times 10^{-15}$ $7.4\times 10^{-4}$
E 0.0338373 1012.39 $4.57\times 10^{-16}$ $3.17\times 10^{-3}$
F 0.145771 1551.5 $1.06\times 10^{-14}$ $1.28\times 10^{-3}$
G 0.232053 965.032 $8.56\times 10^{-14}$ $5.49\times 10^{-4}$
H 0.100534 929.561 $1.12\times 10^{-14}$ $1.34\times 10^{-3}$
I 0.524421 2561.93 $6.78\times 10^{-14}$ $7.27\times 10^{-4}$
Finally, figure 5 presents the total background gravitational wave density (4.1) versus GW frequency, utilizing data from table 4 for selected points.
Figure 5: GW spectrum as a function of backgound gravitational frequency for points in Table 4, considering $v_{w}=1$.
In Figure 5, there are three curves for Lisa, BBO and $\mu$-Ares, which show the power of space-based laser detectors that have not yet been launched into space. As shown in Figure 5, the LISA detector has limited capability to detect the background gravitational waves produced by the electroweak phase transition for the selected points in parameter space. In contrast, the BBO and $\mu$-Ares detectors can detect a portion of these background gravitational waves for all selected points.
The background gravitational wave density from sound wave sources, depicted in all graphs of figure 5, is the primary contributor to the total gravitational wave background at low frequencies. However, as the frequency increases, the contribution from turbulence and collision become dominant. The frequency of approximately $10^{-3}$ Hz marks the maximum values in most plots and delineates the optimal frequency range for detecting the gravitational wave background within our model. As the bubble propagation speed rises from zero to one, the gravitational wave density initially increases, reaching its maximum at around $v_{w}=0.7$, and then decreases thereafter.
In order to understand the effect of changes in the mass of each dark matter component and other independent parameters of the model on gravitational waves, we have changed the values of the parameters around point C and plotted their effects on the values of $\alpha$ and $\frac{\beta}{H_{*}}$ in the graphs of Figures 6 and 7.
Figure 6: Variations in $\alpha$ as a function of changes in each independent parameter around point C in parameter space, with $({\frac{{\Omega}h^{2}}{10^{-15}}})$ represented by color bar.
As shown in Figure 6, increasing the independent dark matter parameters leads to a monotonic rise in $\alpha$ value.
In Figures 6 and 7, we held four parameters constant in each run and varied a single remaining parameter to determine the corresponding gravitational wave parameters.
Figure 7: Variations in $\frac{\beta}{H_{*}}$ as a function of changes in each independent parameter around point C in parameter space, with $T_{p}$ represented by color bar.
According to the graphs in Figure 7, the value of $\frac{\beta}{H_{*}}$ has no significant relationship with the values of dark matter mass, but it has a direct relationship with ${\lambda}_{\phi s}$ and an inverse relationship with ${\nu}_{2}$.
@@ 5 Conclusion
This paper revisits a model of dark matter in which particles exhibit three spin types: a vector particle, a scalar, and a spinor. In our model, the Higgs particle acts as a portal between Standard Model particles and dark matter. We identified some points in parameter space that exceed the detection capability of PandaX-4T and XENONnT, while their relic density matches the Planck satellite’s findings. We examined the electroweak phase transition for selected points and showed that a potential barrier between the two minima of the effective potential at nucleation temperature can generate gravitational waves. The three parameters, $\alpha$, $\frac{\beta}{H_{*}}$, and percolation temperature ($T_{p}$), are sufficient for gravitational wave calculations, independent of other model parameters. Considering the values in Table 4 and the revised background gravitational wave density calculation for each selected point, we find that the ${\Omega h^{2}}_{GW}$ increases with the $\alpha$ value and decreases with the $\frac{\beta}{H_{*}}$ value (i.e., a direct relationship with $\alpha$ and an inverse relationship with $\frac{\beta}{H_{*}}$). Our results indicate that the sensitivity of the LISA detector is insufficient for detecting cosmic background gravitational waves for all benchmarks of Table 2, whereas the BBO and $\mu$-Ares detectors have the capability to confirm the existence of these waves.